# Design and Performance Comparison of FuzzyPID and Non-linear Model Predictive Controller for 4-Wheel Omni-drive Robot

### v1

@@ 7 Conclusions
In this study, we propose a comprehensive framework for developing a trajectory tracking methodology in a 4-wheel Omni-drive robot. This scheme involves deriving the kinematic model, creating an obstacle-free path, followed by a path-smoothing algorithm. Subsequently, we design two popular control methods, namely fuzzyPID and NPMC, for achieving robust path tracking. The entire setup is simulated, and we compare tracking performance by calculating the cross-track error for both controllers.
The results indicate that NPMC outperforms fuzzyPID, albeit with a trade-off in computational complexity. Conversely, the simplicity of fuzzyPID makes it applicable in a wide variety of controller applications. We also assess the effectiveness of NMPC for different values of the prediction horizon in the stability control of our robot.

### v2

@@ 7 Conclusions
In this study, we propose a comprehensive framework for developing a trajectory tracking methodology in a 4-wheel Omni-drive robot. This scheme involves deriving the kinematic model, creating an obstacle-free path, followed by a path-smoothing algorithm. Subsequently, we design two popular control methods, namely fuzzy based PIDs and NPMC, for achieving robust path tracking. The entire setup is simulated, tracking performance is compared by calculating the cross-track error for all of our controllers. Additionally, step-response parameters are computed for clearly highlighting the gap seen in each controllers.
The results indicate that NMPC outperforms fuzzyPIDs, albeit with a trade-off in computational complexity. The NMPC has the better step-response characteristics with low tracking error rate and higher noise rejection capability. Conversely, the simplicity of fuzzyPIDs makes it applicable in a wide variety of controller applications. We also assess the effectiveness of NMPC for different values of the prediction horizon in the stability control of our robot.
@@
8 Limitations and Future Enhancements
While the performance of the aforementioned controllers has been evaluated in simulations, real-time implementation remains untested as they are bound to various external disturbances influencing the stability of mobile robots. Additionally, the simulations were conducted in a static environment limiting their use-case on real world scenarios. Therefore, future work will focus on the real-time implementation of these controllers in both static and dynamic environments to better understand their practical efficacy and adaptability. We also aim to compare the tracking performance of newly proposed IT3 fuzzy system with our proposed approach. Moreover, We also aimed to tune the the sensible gain parameters with the use of various evolutionary algorithms as mentioned in various literature for future work.